# Transfer Learning for Transient Classification: From Simulations to Real Data and ZTF to LSST

### v1

@@ Abstract
Machine learning has become essential for automated classification of astronomical transients, but current approaches face significant limitations: classifiers trained on simulations struggle with real data, models developed for one survey cannot be easily applied to another, and new surveys require prohibitively large amounts of labelled training data. These challenges are particularly pressing as we approach the era of the Vera Rubin Observatory’s Legacy Survey of Space and Time (LSST), where existing classification models will need to be retrained using LSST observations. We demonstrate that transfer learning can overcome these challenges by repurposing existing models trained on either simulations or data from other surveys. Starting with a model trained on simulated Zwicky Transient Facility (ZTF) light curves, we show that transfer learning reduces the amount of labelled real ZTF transients needed by 75% while maintaining equivalent performance to models trained from scratch. Similarly, when adapting ZTF models for LSST simulations, transfer learning achieves 95% of the baseline performance while requiring only 30% of the training data. These findings have significant implications for the early operations of LSST, suggesting that reliable automated classification will be possible soon after the survey begins, rather than waiting months or years to accumulate sufficient training data.
@@
pubyear: 2025pagerange: Transfer Learning for Transient Classification: From Simulations to Real Data and ZTF to LSSTReferences
@@ 1 Introduction
With the development of advanced survey telescopes, we are entering a new era for astronomical study. The Vera Rubin Observatory’s Legacy Survey of Space and Time (LSST) is expected to generate tens of millions of transient alerts per night, an order of magnitude larger than any survey before. This unprecedented volume of data has motivated the development of numerous light-curve classification methods (e.g. Charnock & Moss, 2017; Muthukrishna et al., 2019; Möller & de Boissière, 2020; Boone, 2019; Gomez et al., 2020; Villar et al., 2019, 2020; Carrasco-Davis et al., 2021; Boone, 2021; Qu et al., 2021; Gagliano et al., 2021; Pimentel et al., 2023; Gomez et al., 2023; Sheng et al., 2023; de Soto et al., 2024; Rehemtulla et al., 2024; Shah et al., 2025).
@@
Past studies have been limited by three key constraints. First, classification pipelines are often trained and validated on simulated data, but struggle to generalize to real observations due to subtle differences between simulated and observed light curves (e.g. Muthukrishna et al., 2019). Second, models developed for specific surveys typically require complete retraining when applied to data from different telescopes, despite the underlying physics of the transients remaining unchanged. Third, achieving high classification performance requires large amounts of labelled training data - a resource-intensive process requiring considerable time and expensive spectroscopic follow-up observations.
@@
To address these limitations, we propose using transfer learning, a method that has been successfully applied across many fields in machine learning to adapt models trained for one task to similar applications (Zhuang et al., 2021). The technique works by reusing features and relationships that a model has already learned, even when the specific data distributions differ. This is particularly effective in two scenarios that directly address the challenges in astronomical classification: when labelled data in the target domain are limited or when training new models from scratch is computationally expensive. By reusing learned features and fine-tuning only specific components of the model, transfer learning can dramatically reduce both the amount of labelled data and computational resources needed to achieve high performance on new tasks.
@@
The potential of transfer learning is particularly relevant as we prepare for LSST. It provides a solution to the gap between simulated and real data. Deep learning models require large amounts of training data, which has led to the development of sophisticated light curve simulations (e.g. Kessler et al., 2019a; Narayan & ELAsTiCC Team, 2023) and numerous classifiers trained on these simulations (e.g. Boone, 2019; Hložek et al., 2023). However, simulations do not perfectly capture real observations, making training on labelled survey data necessary (e.g. Rehemtulla et al., 2024). While previous work has treated simulated and real data independently, with models often struggling to generalize between the two, transfer learning provides a framework to leverage features learned from simulations while adapting to real observations. Similarly, while most classifiers are built for specific surveys, transfer learning enables models trained on existing facilities to be adapted for new ones, significantly reducing the data requirements for achieving high performance.
@@
In this work, we demonstrate three key benefits of transfer learning for astronomical transient classification:
@@
    Transfer learning reduces the amount of data needed to generalize from simulated data to real observations
@@
    Transfer learning reduces the amount of labelled data needed to generalize from one survey to another
@@
    Transfer learning reduces training time for classification models, which is important if computing resources are limited
The paper is organized as follows. We introduce the datasets used for our experiments in Section 2. We outline our experimental setup and transfer learning methodology in Section 3. We then examine the benefits of using transfer learning in Section 4, demonstrating both the reduction in required training data and the ability to transfer between simulated and real observations. Finally, we discuss the broader implications of this work and future directions in Section 5.
@@ 2 Data
In this work, we use three distinct datasets to evaluate transfer learning for transient classification: simulated light curves that match the observing properties of the Zwicky Transient Facility (ZTF Sims), real observations from the ZTF Bright Transient Survey (BTS Data), and simulated light curves matching the expected LSST observing properties (LSST Sims). Each dataset provides different opportunities to test transfer learning between simulated and real data, as well as between different surveys. The class distribution for each dataset is summarised in Table 1.
@@
LSST Sims 10424 12272 13000 |——- 8849 (SNIb/c) ——-| |——– 5901 (SNII) ——–| 6014 13000 13000
BTS Data |———- 4454 (SNIa) ———-| |——– 259 (SNIb/c) ——–| |——– 1212 (SNII) ——–| 0 0 0
@@ 2.1 Simulated ZTF Data (ZTF Sims)
The simulated ZTF dataset is described in § 2 of both Gupta et al. (2025) and Muthukrishna et al. (2021), and is based on the models developed for the Photometric LSST Astronomical Time-Series Classification (PLAsTiCC) (Kessler et al., 2019a). Each transient has flux and flux error measurements in the $g$ and $r$ passbands with a median cadence of 3 days in each band [^1] The public MSIP ZTF survey has since changed to a 2-day median cadence. The dataset includes light curves from 12 transient classes: SNIa, SNIa-91bg, SNIax, SNIb, SNIc, SNIc-BL, SNII, SNIIn, SNIIb, TDE, SLSN-I, and AGN.
@@ 2.2 ELAsTiCC (LSST Sims)
The LSST simulations are drawn from the Extended LSST Astronomical Time Series Classification Challenge (ELAsTiCC) dataset (Narayan & ELAsTiCC Team, 2023), which builds upon the original PLAsTiCC simulations with improved models and more realistic survey properties. Light curves are simulated in all six LSST passbands ($ugrizy$) with the expected LSST cadence and depth. We use a subset of classes that overlap with our ZTF simulations: SNIa, SNIax, SNIa-91bg, SNII, SNIb/c, and TDE. Note that some classes are grouped differently than in the ZTF simulations (e.g., SNIb and SNIc are combined).
@@ 2.3 Bright Transient Survey Data (BTS Data)
For real observations, we use data from the ZTF Bright Transient Survey as described in § 2 of Rehemtulla et al. (2024). The BTS systematically classifies every bright ($m<18.5$ mag) extragalactic (and non-AGN) transient discovered by ZTF, providing a spectroscopically complete sample. Light curves are available in the ZTF $g$ and $r$ filters, and we use data from three broad transient classes: SNIa, SNII, and SNIb/c. We use these broader classifications due to the small number of transients classified in most subtypes.
@@ 3.1 Classifiers
We build a Recurrent Neural Network (RNN) classifier with Gated Recurrent Units (GRU; Cho et al., 2014) following the architecture used in Muthukrishna et al. (2019) and updated in Gupta et al. (2024, 2025). As illustrated in Fig. 1, we include a parallel fully-connected (dense) neural network to process any available metadata, which in this work consists only of Milky Way extinction, and we do not use redshift or any host galaxy information. While we present results using this architecture, our transfer learning methodology can work with any neural network design.
The input vector, $\bm{X}_{s}$, for a transient $s$, is a $4\times N_{t}$ matrix, where $N_{t}$ is the maximum number of timesteps for any light curve in our datasets. $N_{t}=656$ in this work, however most light curves have far fewer observations. Each row $i$ of the input matrix $\bm{X}_{s}$ contains the following information:
@@
where $t_{si}$ is the time of the $i$th observation scaled between 0 and 1, $\lambda_{p}$ is the median wavelength of the passband, $f_{si}$ is the flux scaled by dividing by 500, and $\epsilon_{si}$ is the corresponding flux error with the same scaling. We select a constant scaling factor of 500 as it is close to the mean flux in our dataset and enables real-time classification by avoiding the need for future observations. Our model does not require redshift information, making it useful in the early observations of LSST.
@@
where $T$ is the total size of the dataset and $T_{c}$ is the number of transients from class $c$ in the dataset. We train our models using the adam optimizer over $100$ epochs using EarlyStopping implemented in keras.
@@ 3.2 Transfer Learning
Figure 1: Schematic illustrating our neural network architecture and transfer learning methodology. The base model is first trained on ZTF simulations to output probabilities $[P_{b_{1}},P_{b_{2}},...,P_{b_{n}}]$ for each of the $n$ classes in the base dataset. For transfer learning, we replace the output classification layer with a new one that outputs probabilities $[P_{t_{1}},P_{t_{2}},...,P_{t_{m}}]$ for the $m$ classes in the target dataset. When transferring to BTS data, we freeze the base model and only train the new classification layer. When transferring to LSST, we also unfreeze the initial GRU layers to learn the additional passbands while keeping the dense layers frozen. The complexity of the additional Multi-Layer Perceptron (MLP) layers depends on the similarity between the base and target datasets - zero neurons for BTS (which is similar to ZTF sims) and two layers for LSST (which has different passbands).
To assess the effectiveness of transfer learning, we define a base dataset and a target dataset. We first train a classifier on the base dataset, and then adapt it for the target dataset with the specific focus of analysing how much labelled data we need from the target dataset to achieve good performance. For all analyses in this work, ZTF Sims serves as the base dataset.
@@
Fig. 1 summarises our model architecture and experimental setup for transfer learning. We initially train a model to classify a transient into one of the $n$ classes in the ZTF Sims. After pretraining on the ZTF Sims, we replace the classification layer with a Multi-Layer Perceptron (MLP) followed by a new classification layer. The new classification layer varies in size depending on the number of classes, $m$, in the target dataset. The complexity of the MLP is determined by how different the target dataset is from the base dataset. We describe the details of each of these experiments in the following subsections.
@@ 3.2.1 Fine-tuning for BTS
When fine-tuning our model on the Bright Transient Survey, we replace the classification layer with a new one for the BTS classes and freeze all other layers of the base model. Since BTS and ZTF Sims share the same observing properties, we only retrain the new output layer and freeze the rest of the model weights. This prevents the network from modifying the features it has already learned about transients from its pretraining on the ZTF Sims. We use an empty MLP (as seen in Fig. 1), such that there are no intermediate layers connecting the base model to the classification head.
@@ 3.2.2 Fine-tuning for LSST
For LSST, we unfreeze both the initial GRU layers (blue in Fig. 1) and the final classification layers (yellow in Fig. 1). We found that only retraining the final layers, as we did for BTS, resulted in poor performance. The initial GRU layers, which process the time-series information, need to be retrained to effectively handle the more complex and sparser LSST light curves.
An experiment transferring from LSST to ZTF revealed that the choice of which layers to retrain depends on the relationship between the base and target datasets. When going from LSST to ZTF, freezing the entire base model worked best, likely because ZTF’s passbands are effectively a subset of LSST’s. While this direction of transfer is not necessarily useful in practice, it demonstrates the importance of matching the transfer learning approach to the specific datasets involved.
@@ 4.1 Simulated to Real Data
Figure 2: Classification performance as a function of training set size for models trained on BTS data. The orange points show models trained directly on BTS data, while the blue points show models first trained on ZTF Sims and then fine-tuned on BTS data. For transfer learning, we freeze the base model and retrain only the classification layer. Transfer learning achieves equivalent performance to full training while requiring only 25% of the labelled data. The error bars show the standard deviation across 10 independent training runs. The large variance at small training set sizes is due to the limited validation data for determining optimal training duration with Early Stopping. Training on simulations first significantly reduces the amount of data needed to create a good classifier on real data. We use the Area Under the Receiver Operating Characteristics Curve (AUROC) to measure classification performance.
@@
We evaluate the performance using the Area Under the macro-averaged Receiver Operating Characteristics Curve (AUROC). The Receiver Operating Characteristics (ROC) curve plots the False Positive Rate against the True Positive Rate over a range of probability thresholds. We compute the area under the averaged ROC Curve across all classes to quantify classification performance.
Fig. 2 shows that transfer learning from simulated data significantly reduces the amount of real labelled data needed to build an effective classifier. A model pretrained on ZTF Sims requires only 25% of the BTS data to match the performance of a model trained from scratch, and achieves 90% of the baseline performance with just 5% of the data. This reduction in required training data is particularly valuable given the time and expense of obtaining spectroscopic labels.
The large variance in performance at small training set sizes (see Fig. 2) is due to the limited validation data available for determining optimal training duration with Early Stopping. However, even with this uncertainty, transfer learning consistently outperforms training from scratch when labelled data is limited.

### v2

@@ Abstract
Machine learning has become essential for automated classification of astronomical transients, but current approaches face significant limitations: classifiers trained on simulations struggle with real data, models developed for one survey cannot be easily applied to another, and new surveys require prohibitively large amounts of labelled training data. These challenges are particularly pressing as we approach the era of the Vera C. Rubin Observatory’s Legacy Survey of Space and Time (LSST), where existing classification models will need to be retrained using LSST observations. We demonstrate that transfer learning can overcome these challenges by repurposing existing models trained on either simulations or data from other surveys. Starting with a model trained on simulated Zwicky Transient Facility (ZTF) light curves, we show that transfer learning reduces the amount of labelled real ZTF transients needed by 95% while maintaining equivalent performance to models trained from scratch. Similarly, when adapting ZTF models for LSST simulations, transfer learning achieves 94% of the baseline performance while requiring only 30% of the training data. These findings have significant implications for the early operations of LSST, suggesting that reliable automated classification will be possible soon after the survey begins, rather than waiting months or years to accumulate sufficient training data.
@@
pubyear: 2025pagerange: Transfer Learning for Transient Classification: From Simulations to Real Data and ZTF to LSSTB
@@ 1 Introduction
With the development of advanced survey telescopes, we are entering a new era for astronomical study. The Vera Rubin Observatory’s Legacy Survey of Space and Time (LSST; Ivezić et al., 2019) is expected to generate tens of millions of transient alerts per night, an order of magnitude more than current state-of-the-art surveys. This unprecedented volume of data has motivated the development of numerous machine learning-based alert and light-curve classification methods (e.g. Charnock & Moss, 2017; Muthukrishna et al., 2019; Möller & de Boissière, 2020; Boone, 2019; Gomez et al., 2020; Villar et al., 2019, 2020; Carrasco-Davis et al., 2021; Boone, 2021; Qu et al., 2021; Gagliano et al., 2021; Muthukrishna et al., 2022; Pimentel et al., 2023; Gomez et al., 2023; Sheng et al., 2024; de Soto et al., 2024; Rehemtulla et al., 2024; Shah et al., 2025).
@@
Past studies have been limited by three key constraints. First, classification pipelines are often trained and validated on simulated data, but struggle to generalise to real observations due to subtle differences between simulated and observed light curves (e.g., Muthukrishna et al., 2019). Second, models developed for specific surveys typically require complete retraining when applied to data from different telescopes, despite the underlying physics of the transients remaining unchanged. Third, achieving high classification performance requires large amounts of labelled training data - a resource-intensive process requiring considerable time and expensive spectroscopic follow-up observations.
@@
To address these limitations, we explore using transfer learning, a method that has been successfully applied across many fields in machine learning to adapt models trained for one task to similar applications (Zhuang et al., 2021). The technique works by reusing features and relationships that a model has already learned, even when the specific data distributions differ. This is particularly effective in two scenarios that directly address the challenges in astronomical classification: when labelled data in the target domain are limited or when training new models from scratch is computationally expensive. By reusing learned features and fine-tuning only specific components of the model, transfer learning can dramatically reduce both the amount of labelled data and computational resources needed to achieve high performance on new tasks.
@@
The potential of transfer learning and domain adaptation is particularly relevant as we prepare for LSST. It provides a solution to the gap between simulated and real data. Deep learning models require large amounts of training data, which has led to the development of sophisticated light curve simulations (e.g., Kessler et al., 2019; Narayan & ELAsTiCC Team, 2023) and numerous classifiers trained on these simulations (e.g., Hložek et al., 2023). However, simulations do not perfectly capture real observations, making training on labelled survey data necessary (e.g., Rehemtulla et al., 2024). While previous work has treated simulated and real data independently, with models often struggling to generalise between the two, transfer learning provides a framework to leverage features learned from simulations while adapting to real observations. Similarly, while most classifiers are built for specific surveys, transfer learning enables models trained on existing facilities to be adapted for new ones, significantly reducing the data requirements for achieving high performance. However, the initial model architecture determines the feasibility of transfer learning across different surveys, and not all existing classification models are suitable candidates for this approach. In this work, we incorporate the observation wavelength directly into the model to facilitate transfer learning across different surveys with varying passbands.
@@
We demonstrate three key benefits of transfer learning for astronomical transient classification:
@@
    Transfer learning reduces the amount of data needed to generalise from simulated data to real observations,
@@
    Transfer learning reduces the amount of labelled data needed to generalise from one survey to another, and
@@
    Transfer learning reduces training time for classification models.
The paper is organised as follows. We introduce the datasets used for our experiments in Section 2. We outline our experimental setup and transfer learning methodology in Section 3. We then examine the benefits of using transfer learning in Section 4, demonstrating both the reduction in required training data and the ability to transfer between simulated and real observations. Finally, we discuss the broader implications of this work and future directions in Section 5.
@@ 2 Data
In this work, we use three distinct datasets to evaluate transfer learning for transient classification: simulated light curves that match the observing properties of the Zwicky Transient Facility (ZTF; Bellm et al., 2019), real observations from the ZTF Bright Transient Survey (BTS; Fremling et al., 2020; Perley et al., 2020; Rehemtulla et al., 2024), and simulated light curves matching the expected LSST observing properties. Each dataset provides different opportunities to test transfer learning between simulated and real data, as well as between different surveys. The class distribution for each dataset is summarised in Table 1.
@@
LSST Sims 51745 12272 12012 37261 24513 10331 71806 34016 23408 28285 28322 32681
BTS Data |———- 4454 (SNIa) ———-| |——– 259 (SNIb/c) ——–| |——– 1212 (SNII) ——–| 22 46 0
@@ 2.1 Simulated ZTF Data (ZTF Sims)
The simulated ZTF dataset is described in §2 of both Gupta et al. (2025) and Muthukrishna et al. (2022), and is based on the models developed for the Photometric LSST Astronomical Time-Series Classification (PLAsTiCC; Kessler et al., 2019). Each transient has flux and flux error measurements in the $g$ and $r$ passbands with a median cadence of 3 days in each band [^1] The public MSIP ZTF survey has since changed to a 2-day median cadence.. The dataset includes light curves from 12 transient classes: SNIa, SNIa-91bg, SNIax, SNIb, SNIc, SNIc-BL, SNII, SNIIn, SNIIb, TDE, SLSN-I, and AGN.
@@ 2.2 ELAsTiCC (LSST Sims)
The LSST simulations are drawn from the Extended LSST Astronomical Time Series Classification Challenge (ELAsTiCC; Narayan & ELAsTiCC Team, 2023), which builds upon the original PLAsTiCC simulations with improved models and more realistic survey properties. Light curves are simulated in all six LSST passbands ($ugrizy$) with the expected LSST cadence and depth. ELAsTiCC provides realistic class distributions based on expected LSST observation rates [^2] ELAsTiCC also includes upsampled rare classes such as kilonovae, but we do not use these in this work.. We use the same 12 transient classes as our ZTF simulations, with class distributions shown in Table 1 reflecting the expected rates for LSST observations.
@@ 2.3 Bright Transient Survey Data (BTS Data)
For real observations, we use data from the ZTF BTS. The aim of BTS is to systematically classify every bright ($m<18.5$ mag) extragalactic transient discovered in the public ZTF alert stream (Masci et al., 2019), providing a spectroscopically complete sample. The adopted dataset is queried and processed following the procedures in Rehemtulla et al. (2024) and considers data up to August 2024. Light curves are available in the ZTF $g$ and $r$ filters, and we use data from three broad transient classes: SNIa, SNII, and SNIb/c, and two rarer transient classes, TDE and SLSN-I. We use these broader classifications because of the small number of transients classified in most subtypes.
@@ 3.1 Classifiers
We build a Recurrent Neural Network (RNN) classifier with Gated Recurrent Units (GRU; Cho et al., 2014) following the architecture used in Muthukrishna et al. (2019) and updated in Gupta et al. (2024, 2025). As illustrated in Figure 1, we include a parallel fully-connected (dense) neural network to process any available metadata, which in this work consists only of Milky Way extinction, and we do not use redshift or any host galaxy information. Contextual information such as host galaxy properties have been shown to be very useful for transient classification (Foley & Mandel, 2013; Gagliano et al., 2021) and while we do not use host information in our work, our classification framework is designed to incorporate both time-dependent and time-independent features.
While we present results using this architecture, we expect transfer learning results to be broadly consistent across other classification architectures that use flexible input formats capable of handling data from different surveys.
To enable transfer learning in this work, we use a survey-agnostic input representation. The input vector, $\bm{X}_{s}$, for a transient $s$, is a $4\times N_{t}$ matrix, where $N_{t}$ is the maximum number of timesteps for any light curve in our datasets. $N_{t}=656$ in this work, however most light curves have far fewer observations. Each row $i$ of the input matrix $\bm{X}_{s}$ contains the following information:
@@
where $t_{si}$ is the time of the $i$th observation scaled between 0 and 1, $\lambda_{p}$ is the median wavelength of the passband, $f_{si}$ is the flux scaled by dividing by 500, and $\epsilon_{si}$ is the corresponding flux error with the same scaling. We select a constant scaling factor of 500 as it is close to the mean flux in our dataset and enables real-time classification by avoiding peak-based scaling, which would depend on observations that may not yet be available in real-time scenarios. Our model does not require redshift information, making it useful in the early observations of LSST when redshifts may not be known for many sources.
@@
where $T$ is the total size of the dataset and $T_{c}$ is the number of transients from class $c$ in the dataset. We train our models using the Adam optimiser (Kingma & Ba, 2014) over $100$ epochs using EarlyStopping implemented in keras (Chollet, 2015).
@@ 3.2 Transfer Learning
Figure 1: Schematic illustrating our neural network architecture and transfer learning methodology. The model is first pretrained on ZTF simulations to output probabilities $[P_{b_{1}},P_{b_{2}},...,P_{b_{n}}]$ for each of the $n$ classes in the source dataset. For transfer learning, we replace the output classification layer with a new one that outputs probabilities $[P_{t_{1}},P_{t_{2}},...,P_{t_{m}}]$ for the $m$ classes in the target dataset. When transferring to BTS data, we freeze the pretrained model and only train the new classification layer. When transferring to LSST, we also unfreeze the initial GRU layers to learn the additional passbands while keeping the dense layers frozen. The complexity of the additional Multi-Layer Perceptron (MLP) layers depends on the similarity between the source and target datasets - zero neurons for BTS (which is similar to ZTF sims) and two layers for LSST (which has different passbands).
To assess the effectiveness of transfer learning, we define a source dataset and a target dataset. We first train a classifier on the source dataset, and then adapt it for the target dataset with the specific focus of analysing how much labelled data we need from the target dataset to achieve good performance. For all analyses in this work, ZTF Sims serves as the source dataset.
@@
Figure 1 summarises our model architecture and experimental setup for transfer learning. We initially train a model to classify a transient into one of the $n$ classes in the ZTF Sims. After pretraining on the ZTF Sims, we replace the classification layer with a Multi-Layer Perceptron (MLP) followed by a new classification layer. The new classification layer varies in size depending on the number of classes, $m$, in the target dataset. The complexity of the MLP is determined by how different the target dataset is from the source dataset. We describe the details of each of these experiments in the following subsections.
We note that the primary benefit of transfer learning comes from pretraining rather than from specific choices about which layers to freeze, as unfreezing the entire network yields similar performance. The choice of freezing strategy primarily acts to reduce computational costs during fine-tuning.
@@ 3.2.1 Fine-tuning for BTS
When fine-tuning our model on the BTS data, we replace the classification layer with a new one for the $m=5$ BTS classes and freeze all other layers of the pretrained model. Since BTS and ZTF Sims share the same observing properties, we only retrain the new output layer and freeze the rest of the model weights. This prevents the network from modifying the features it has already learned about transients from its pretraining on the ZTF Sims.
@@ 3.2.2 Fine-tuning for LSST
For LSST, we unfreeze both the initial GRU layers (blue in Figure 1) and the final classification layers (yellow in Figure 1). We found that only retraining the final layers, as we did for BTS, resulted in poor performance. The initial GRU layers, which process the time-series information, need to be retrained to effectively handle the more complex and sparser LSST light curves.
While unfreezing the entire network performs similarly to our selective freezing approach, we choose to freeze the dense layers to reduce computational costs while maintaining equivalent performance.
@@ 4 Results
Refer to caption Refer to caption
Figure 2: Transfer learning performance from ZTF sims to BTS [left] and ZTF Sims to LSST Sims [right]. The x-axis shows the number of randomly sampled training light curves provided to the model and the y-axis shows the classification performance as the macro-averaged Area Under the Receiver Operating Characteristic (ROC) curve for each class (AUROC). The lines and shaded regions show the mean and standard deviation across ten independent training runs, respectively.
@@ 4.1 Simulated to Real Data
We evaluate the performance using the Area Under the macro-averaged Receiver Operating Characteristics Curve (AUROC). The Receiver Operating Characteristic (ROC) curve plots the False Positive Rate against the True Positive Rate over a range of classification probability thresholds.
Figure 2 [left] shows that transfer learning from simulated data significantly reduces the amount of real labelled data needed to build an effective classifier. A model pretrained on ZTF Sims achieves the baseline performance with just 5% of the data. This reduction in required training data is particularly valuable given the time and expense of obtaining spectroscopic labels. We restrict our evaluation at 4,000 labelled training samples as going further would leave too few TDEs in the test set for robust evaluation. Even at this point, transfer learning provides a significant performance improvement when compared to direct training.
The performance improvement in low-data scenarios is primarily driven by minority classes, particularly TDEs and SLSNe-I, which transfer learning can effectively classify despite having very few labelled examples. These rare classes also account for the large variance observed at small training set sizes in Figure 2, where classifiers become highly sensitive to each additional training sample. The class-specific results in Figure 3 confirm that most performance variability stems from the TDE and SLSN-I classes. Importantly, transfer learning still provides substantial benefits when applied to only the common transient classes, as demonstrated in Appendix B.